# MMiC: Mitigating Modality Incompleteness in Clustered Federated Learning

### v1

@@ 4 Methodology
Figure 1: The MMiC architecture is divided into three layers: Server, Cluster, and Client. The right part illustrates the process of parameter substitution within a cluster. The MPO module primarily operates during the aggregation process from clusters to the global model, while the Banzhaf client selection is applied to the client selection within each cluster.
@@
Each cluster accommodates multiple clients, and each client hosts multimodal data with missing modalities locally. The cluster model is derived from the federated average aggregation (Section 3.1). Within each cluster, MMiC handles the missing modality issue and the client selection task. First, a parameter layer substitution mechanism allows clients with missing modalities to obtain knowledge from clients with complete modality data within the same cluster (Section 4.1). Second, client selection is performed leveraging the Banzhaf Power Index, which identifies a set of core clients that contribute the most effectively to the cluster’s model (Section 4.2).
For global aggregation, MMiC extends FedAdagrad by adopting the Markowitz Portfolio Optimization (Section 4.3). This method incorporates client training evaluation results from each cluster and reduces the impact of missing modalities by down-weighing the contribution from clusters that contain missing modalities.
4.1 Cluster-based Parameter Substituion
We are targeting the multimodal scenario when the dataset has both textual $X^{text}$ and image $X^{image}$ data, thus we form the input as a pair {$X^{text}_{i},X^{image}_{i}$} to the $i$-th data in $\mathcal{D}$. The missing modality randomly occurs resulting in data samples in the form of {$X^{null}_{i},X^{image}_{i}$} or {$X^{text}_{i},X^{null}_{i}$} [^2] Please note that MMiC is modality-agnostic – it remains effective regardless of the type or number of modalities, as long as modalities are consistent across clients. We have additional experiments on other modalities shown in Section 5.6.
@@ 4.1 Cluster-based Parameter Substituion
Poverty Parameter Identification. In one round of client model training, we identify the layer that is the most affected by missing modalities as $poverty$ layer. We use the parameter difference between the last batch $b-1$ and the current batch $b$ to define the degree of this effect when the missing modality occurs in batch $b$ and batch $b-1$ is with complete modality [^3] We are aware of the missing modality status of the data.:
$\Delta\boldsymbol{\theta}_{k,t}^{e,b}(p)$ $=\text{mean}\left(\left|\frac{\boldsymbol{\theta}_{k,t}^{e,b}(p)-\boldsymbol{\theta}_{k,t}^{e,b-1}(p)}{\boldsymbol{\theta}_{k,t}^{e,b-1}(p)}\right|\right)$
$\text{where}\quad\Delta\boldsymbol{\theta}_{k,t}^{e,b}(p)\in\mathbb{R}$ (8)
where $e$ is the epoch, $t$ is the training round, $k$ is the client. $\Delta\boldsymbol{\theta}_{k,t}^{e,b}(p)$ represents average parameters change rate of the $p$-th layer. As $\Delta\boldsymbol{\theta}_{k,t}^{e,b}$ is a set that contains all the layers change rate. We collect all the batches in the $e$-th epoch that contain complete modality data, while the preceding batch contains missing modality data, and form a batch set $\mathcal{B}*$. Therefore, the relative change rate for the layer $p$ in one epoch could be expressed as:
$\Delta\boldsymbol{\theta}_{k,t}^{e}(p)=\frac{\sum_{b\in\mathcal{B}^{*}}\Delta\boldsymbol{\theta}_{k,t}^{e,b}(p)}{|\mathcal{B}^{*}|}$ (9)
After calculating the average rate of change for all layers in the model in one epoch, we identify the layer with the highest rate of change across all epochs. This layer would be considered as the $poverty$ layer with index $\mathcal{P}_{k,t}$:
$\mathcal{P}_{k,t}=\arg\max_{\begin{subarray}{c}p\in\{1,2,\dots,P\}\end{subarray}}\Delta\boldsymbol{\theta}_{k,t}(p)$ (10)
@@
Poverty Parameter Substitution (PPS). After identifying the poverty parameter in $c_{k}$, we propose to replace these parameters with the parameters from other clients in the cluster that have complete modality data. Specifically, we compute a weighted average of the corresponding parameters of other clients and replace the poverty parameter in $c_{k}$.
$\boldsymbol{\theta}_{k,t}(\mathcal{P}_{k,t})$ $=\sum_{j=1}^{J}w_{j}\boldsymbol{\theta}_{j,t}(\mathcal{P}_{k,t})$
$\quad\text{where }\mathcal{P}_{j,t}=\varnothing,\quad K+J=|\mathcal{M}_{m}|$ (11)
$c_{k}$ represents the poverty client and $c_{j}$ is the client that trained with the complete modality round, and both $c_{k}$ and $c_{j}$ belong to the cluster $\mathcal{M}_{m}$.
Optimizers Could Approximate Parameter Difference. Saving parameters between two batches is resource-intensive and time-consuming. However, we can roughly deduce the rate of change by utilizing the gradient derivation function of the optimizer. Below, we provide the calculation process for parameter variations during the SGD and Adam optimization processes. For SGD, $\eta$ is the learning rate and $\boldsymbol{g}_{k,b}$ is the gradient for $c_{k}$ in $b$ batch.
$\Delta\boldsymbol{\theta}_{k,t}^{e,b}=\frac{(\boldsymbol{\theta}_{k,t}^{e,b-1}-\eta_{k,t}\boldsymbol{g}_{k,b})-\boldsymbol{\theta}_{k,t}^{e,b-1}}{\boldsymbol{\theta}_{k,t}^{e,b-1}}\\ =\frac{-\eta_{k,t}\boldsymbol{g}_{k,b}}{\boldsymbol{\theta}_{k,t}^{e,b}+\boldsymbol{g}_{k,b}}$ (12)
Similarly, we can utilize the Adam optimizer to deduce the average rate of change, here $\hat{\boldsymbol{momentum}}$ is the bias-corrected moment estimates and $\hat{\boldsymbol{\nu}}$ is the second-moment estimate in Adam optimizer (Kingma & Ba, 2015):
$\Delta\boldsymbol{\theta}_{k,t}^{e,b}=\frac{\eta_{k,t}abs(\-\hat{\boldsymbol{momentum}}_{b})}{abs(\boldsymbol{\theta}_{k,t}^{e,b})(\sqrt{\hat{\boldsymbol{\nu}}_{b}}+\epsilon)}$ (13)
4.2 Banzhaf Client Selection in Clusters
@@
Few CFL work addresses this issue and quantifies the impact of client selection for each cluster. We propose a novel method to fill the gap by leveraging Banzhaf Power Index (BPI) to select the core clients. It comprises two main steps. First, we identify core members within each subset selected in each round and tally the times of core members, denoted as core times. Then, by leveraging the historical selection times and core times, we devise a new client selection probability formula, which is applied in our subsequent client selection process. Banzhaf client selection does not require extra communication.
Core Members Identification. At training round $t$, each cluster $\mathcal{M}_{m}$ makes a random client selection. We can obtain the $i$-th client’s local model performance $a_{i,t}$ (i.e., accuracy, F1 score or RSum) of each client in the $t$-th round, and the performance $a_{i,t-1}$ in the ($t$-1)-th round. We calculate the return $\alpha_{i,t}$ of the client for the $t$-th round based on the difference between these two rounds of performances as follows:
$\alpha_{i,t}=a_{i,t}-a_{i,t-1}$ (14)
@@
$A_{\mathcal{S}_{m,t}}=\sum_{i=0}^{|\mathcal{S}_{m,t}|}w_{i}\alpha_{i,t}$ (15)
For each cluster $\mathcal{S}_{m,t}$, we have its total return $A_{\mathcal{S}_{m,t}}$ by Equation (15). Correspondingly, we can calculate the total return for $\mathcal{S}_{m,t}^{{}^{\prime}}$ when excluding client $c_{i}$, $\𝑖\mathcal{S}_{m,t}^{\prime}=\mathcal{S}_{m,t}\textbackslash\{i\}$: Then, we use the return of cluster model to evaluate the core members:
$c_{i}\,is\,core\,member?\begin{cases}True&A_{\mathcal{S}_{m,t}}\geq\alpha^{m}_{t}\text{ and }A_{\mathcal{S}_{m,t}^{{}^{\prime}}}<\alpha^{m}_{t}\\ False&\text{otherwise}\end{cases}$ (16)
@@
$\phi(i)=\sum_{t=1}^{T}\mathbb{I}\left(c_{i}\,is\,a\,core\,member\,\right)$ (17)
@@
Client Selection Probability. We could design a method to measure the probability of a client being selected in the current round based on its core times. Our goal is to ensure that a higher core time corresponds to a higher probability. However, this way will make clients with small core times unlikely to be selected later. To solve this problem, we set another counter $\mathcal{T}(*)$ that records the number of times the client has been selected in its history, and we determine the final selection probability by balancing the historically selected times $\mathcal{T}(i)$ and the core times $\phi(i)$:
$prob(i)=\frac{e^{\frac{\tau\phi(i)}{\mathcal{T}(i)}}}{\sum_{i\in\mathcal{M}_{m}}e^{\frac{\tau\phi(i)}{\mathcal{T}(i)}}}$ (18)
@@
4.3 Markowitz Profolio Optimization for Global Aggregation
As introduced in Section 3.3, Markowitz Portfolio Optimization is a framework for constructing an investment portfolio that aims to maximize returns for a given level of risk. We are the first to integrate this theory into CFL by considering to mitigate the negative effects of local client models (e.g., those with missing modalities) on the global model. Specifically, we propose a mechanism where, in each communication round, the performance of the selected clients within a cluster $\mathcal{M}_{m}$ is quantified through a non-linear scoring function. This score serves as a direct indicator of the cluster’s performance during the round, allowing for adaptive control over the magnitude of the global model’s updates. By leveraging this approach, we dynamically regulate the contribution of each cluster to the global model aggregation, thereby reducing the impact of underperforming or unreliable clients while maintaining the global model’s robustness.
To achieve this, we require three core indicators to construct our “investment portfolio” for the cluster $\mathcal{M}m$: the return (i.e., payoff) of each cluster ($A_{m,t}$, Equation (15)); the risk associated with the cluster ($\bar{\sigma}_{m,t}$); and the tolerance for return-risk ratio for each cluster ($\lambda$).
@@ 4.3 Markowitz Profolio Optimization for Global Aggregation
Cluster Risk. According to the asset risk calculation formula Equation (7), we obtain the covariance as follows:
$\bar{\sigma}_{m,t}$ $=\sum_{i=1}^{|\mathcal{S}_{m}|}\sum_{j=1}^{|\mathcal{S}_{m}|}w_{i}w_{j}\Bigg{\{}\frac{1}{T-1}\sum_{t=1}^{T}$
$\quad(\alpha_{i,t}-\bar{\alpha}_{i,t-1})(\alpha_{j,t}-\bar{\alpha}_{j,t-1})\Bigg{\}}$ (19)
@@
where $\bar{\alpha}_{i,t-1}$ represents the historical average return for client $i$ in round $t-1$. $w_{i}$ and $w_{j}$ are weights assigned to the $i$-th and $j$-th clients respectively. Please refer to Section 3.1 for weights’ calculation. Then the cluster risk is formulated as follows according to Equation (7):
$\quad\gamma_{m,t}=\lambda\bar{\sigma}_{m,t}-(1-\lambda)A_{m,t}$ (20)
where $\lambda$ [0,1] is the risk tolerance constant. We defaultly set it as 0.5 to balance risk and return. Our goal is to minimize the sum of weighted cluster RARC (cluster weight calculation is introduced in Section 3.1):
$minimize\quad\gamma_{t}=\sum_{m=1}^{M}w_{m}\gamma_{m,t}$ (21)
We analyzed how we achieved the time complexity of MPO in Appendix D.
@@
Participate in Global Aggregation. A higher $\gamma_{t}$ signifies that the clusters have undergone a relatively poor local client training process. In such scenarios, our objective is to diminish the impact of the current round’s parameter updates on the global aggregation by down-weighting the cluster parameters associated with higher $\gamma$. To accomplish this, we introduce a dynamic $\beta^{*}$ to replace the static $\beta$ in FedAdagrad, resulting in FedAdagradMPO, which effectively captures the influence of $\gamma_{t}$:
$\beta^{*}=\beta+\frac{tanh(ReLU(\gamma_{t}))}{\tau}$ (22)
@@
When $\gamma_{t}>0$, then $\beta^{*}>\beta$, which helps retain more historical parameter updates in the global aggregation. Otherwise, $\beta^{*}=\beta$. Here $\tau$ serves as the smoothness control constant for the range. We suggest setting the temperature to range [10,20], and it depends on the value set for $\beta$.
4.4 Variants of MMiC

### v3

@@ 4. Methodology
Figure 1. The MMiC architecture is divided into three layers: Server, Cluster, and Client. The right part illustrates the process of parameter substitution within a cluster. The MPO module primarily operates during the aggregation process from clusters to the global model, while the Banzhaf client selection is applied to the client selection within each cluster.
[…]
@@
Each cluster accommodates multiple clients, and each client hosts multimodal data with missing modalities locally. The cluster model is derived from the federated average aggregation (Section 3). Within each cluster, MMiC handles the missing modality issue and the client selection task. First, a parameter layer substitution mechanism allows clients with missing modalities to obtain knowledge from clients with complete modality data within the same cluster (Section 4.1). Second, client selection is performed leveraging the Banzhaf Power Index, which identifies a set of core clients that contribute the most effectively to the cluster’s model (Section 4.2).
For global aggregation, MMiC extends $FedAdagrad$ (Reddi et al., 2021a) by adopting the Markowitz Portfolio Optimization (Section 4.3). This method incorporates client training evaluation results from each cluster and reduces the impact of missing modalities by down-weighting the contribution from clusters that contain missing modalities.
4.1. Cluster-based Parameter Substituion
We consider a multimodal setting where $i$-th data instance is a pair $\{X^{text}_{i},X^{image}_{i}\}$. Missing modalities occur randomly, yielding samples such as $\{X^{null}_{i},X^{image}_{i}\}$ or $\{X^{text}_{i},X^{null}_{i}\}$ [^1] Please note that MMiC is modality-agnostic – it remains effective regardless of the type or number of modalities, as long as modalities are consistent across clients. We have additional experiments on other modalities shown in Section 5.4.
@@ 4.1. Cluster-based Parameter Substituion
Poverty Parameter Identification. In one round of client model training, we identify the layer that is the most affected by missing modalities as povertypoverty layer. We use the parameter difference between the last batch bb-1 and the current batch $b$ to define the degree of this effect when the missing modality occurs in batch $b$ and batch b1b-1 is with complete modality [^2] We are aware of the missing modality status of the data.:
(2) $\rho_{k,t}^{e,b}(p)=mean\left(\left|\frac{\bm{\theta}_{k,t}^{e,b}(p)-\bm{\theta}_{k,t}^{e,b-1}(p)}{\bm{\theta}_{k,t}^{e,b-1}(p)}\right|\right)\;\text{and}\;\rho_{k,t}^{e,b}(p)\in\mathbb{R}$
where $e$ represents the epoch, $t$ is the training round, $k$ denote the client. $\rho_{k,t}^{e,b}(p)$ represents average parameter change rate of the $p$-th layer. As $\rho_{k,t}^{e,b}$ is a set containing all layers’ change rates. We collect all the batches in the $e$-th epoch that contain complete modality data, while the preceding batch contains missing modality data, and form a batch set $\mathcal{B}*$. Therefore, the relative change rate for the layer $p$ in one epoch could be expressed as:
(3) $\rho_{k,t}^{e}(p)=\frac{\sum_{b\in\mathcal{B}^{*}}\rho_{k,t}^{e,b}(p)}{|\mathcal{B}^{*}|}$
After calculating the average change rate for all layers in the model in one epoch, we identify the layer with the highest change rate across all epochs. This layer would be considered as the povertypoverty layer with index $\mathcal{P}_{k,t}$:
(4) $\mathcal{P}_{k,t}=\arg\max_{\begin{subarray}{c}p\in\{1,2,\dots,P\}\end{subarray}}\rho_{k,t}(p)$
@@
Poverty Parameter Substitution (PPS). After identifying the poverty parameter in $c_{k}$, we propose to replace these parameters with the parameters from other clients in the cluster that have complete modality data. The proposed parameter substitution can be considered as a model editing strategy (Li et al., 2024). Specifically, we compute a weighted average of the corresponding parameters of other clients and replace the poverty parameter in $c_{k}$.
(5) $\bm{\theta}_{k,t}(\mathcal{P}_{k,t})=\sum_{j=1}^{J}{w_{j}}\bm{\theta}_{j,t}(\mathcal{P}_{k,t})\quad where\quad\mathcal{P}_{j,t}=\varnothing,K+J=|\mathcal{M}_{m}|$
where $c_{k}$ represents the poverty client and $c_{j}$ is the client that trained with the complete modality round, and both $c_{k}$ and $c_{j}$ belong to the cluster $\mathcal{M}_{m}$.
Optimizers Could Approximate Parameter Difference. Saving parameters between two batches is resource-intensive and time-consuming. However, we can roughly deduce the rate of change by utilizing the gradient derivation function of the optimizer. Below, we provide the calculation process for parameter variations during the Adam optimization processes. Here $\hat{\bm{momentum}}$ is the bias-corrected moment estimate and $\hat{\bm{\nu}}$ is the second-moment estimate in Adam optimizer (Kingma and Ba, 2015):
(6) $\Delta\bm{\theta}_{k,t}^{e,b}=\frac{\eta_{k,t}abs(\-\hat{\bm{momentum}}_{b})}{abs(\bm{\theta}_{k,t}^{e,b})(\sqrt{\hat{\bm{\nu}}_{b}}+\epsilon)}$
4.2. Banzhaf Client Selection in Clusters
@@
Few CFL work addresses this issue and quantifies the impact of client selection for each cluster. We propose a novel method to fill the gap by leveraging Banzhaf Power Index (Dubey and Shapley, 1979; Laruelle and Valenciano, 2001) (BPI) to select the core clients. It comprises two main steps. First, we identify core members within each subset selected in each round and tally the times of core members, denoted as core times. Then, by leveraging the historical selection times and core times, we devise a new client selection probability formula, which is applied in our subsequent client selection process. Banzhaf client selection does not require extra communication.
Core Members Identification. At training round $t$, each cluster $\mathcal{M}_{m}$ makes a random client selection. We can obtain the $i$-th client’s local model performance $a_{i,t}$ (i.e., accuracy, F1 score or RSum) of each client in the $t$-th round, and the performance $a_{i,t-1}$ in the (tt-1)-th round. We calculate the return $\alpha_{i,t}$ of the client for the $t$-th round based on the difference between these two rounds of performances as follows:
(7) $\alpha_{i,t}=a_{i,t}-a_{i,t-1}$
@@
(8) $A_{\mathcal{S}_{m,t}}=\sum_{i=0}^{|\mathcal{S}_{m,t}|}w_{i}\alpha_{i,t}$
[…]
(9) $c_{i}\,is\,core\,member?\begin{cases}True&A_{\mathcal{S}_{m,t}}\geq\alpha^{m}_{t}\text{ and }A_{\mathcal{S}_{m,t}^{{}^{\prime}}}<\alpha^{m}_{t}\\ False&\text{otherwise}\end{cases}$
@@
(10) $\phi(i)=\sum_{t=1}^{T}\mathbb{I}\left(c_{i}\,is\,a\,core\,member\,\right)$
@@
Client Selection Probability. BPI prioritizes clients without missing modalities, thus favoring those that contribute more to performance improvements. We could design a method to measure the probability of a client being selected in the current round based on its core times. Our goal is to ensure that a higher core time corresponds to a higher probability. However, this way will make clients with small core times unlikely to be selected later. To solve this problem, we set another counter $\mathcal{T}(*)$ that records the number of times the client has been selected in its history, and we determine the final selection probability by balancing the historically selected times $\mathcal{T}(i)$ and the core times $\phi(i)$:
(11) $prob(i)$ $=\frac{e^{\frac{\tau\phi(i)}{\mathcal{T}(i)}}}{\sum_{i\in\mathcal{M}_{m}}e^{\frac{\tau\phi(i)}{\mathcal{T}(i)}}}$
@@
4.3. Markowitz Portfolio Optimization for Global Aggregation
In our framework, MMiC builds upon $FedAdagrad$ (Reddi et al., 2021a) to support adaptive cluster-level aggregation in the presence of modality incompleteness. To enhance global robustness, we propose a novel integration of Markowitz Portfolio Optimization (MPO) into $FedAdagrad$.
Originally introduced in financial theory (Miller, 1960), MPO constructs investment portfolios that trade off expected return and associated risk. Inspired by this principle, we adapt the concept to dynamically regulate the influence of each cluster in global aggregation. Specifically, clusters consisting of underperforming clients—such as those suffering from missing modalities—may introduce biased or unreliable updates. We treat each cluster as an “investment” and aim to down-weight risky clusters while preserving beneficial contributions.
Our approach introduces a scoring function that quantifies the overall quality of a cluster’s update in each round, based on both its expected return and internal risk. This score directly informs the global aggregation process, enabling us to balance between current update quality and accumulated historical knowledge. To achieve this, we require three core indicators to construct our “investment portfolio” for cluster $\mathcal{M}m$ includes the expected return (i.e., payoff) of the cluster, denoted as $A_{m,t}$ (Equation 8); the risk $\bar{\sigma}_{m,t}$, derived from the intra-cluster covariance; and the return-risk tolerance coefficient $\lambda$ that governs the trade-off between return and risk for the cluster.
@@ 4.3. Markowitz Portfolio Optimization for Global Aggregation
Cluster Risk. According to the asset risk calculation formula from MPO (Miller, 1960; Markowitz, 1991), we obtain the covariance as follows:
(12) $\bar{\sigma}_{m,t}=\sum_{i=1}^{|\mathcal{S}_{m}|}\sum_{j=1}^{|\mathcal{S}_{m}|}w_{i}w_{j}\left\{\frac{1}{T-1}\sum_{t=1}^{T}(\alpha_{i,t}-\bar{\alpha}_{i,t-1})(\alpha_{j,t}-\bar{\alpha}_{j,t-1})\right\}$
@@
where $\bar{\alpha}_{i,t-1}$ represents the historical average return for client $i$ in round t1t-1. $w_{i}$ and $w_{j}$ are weights assigned to the $i$-th and $j$-th clients respectively (Please refer to Section 3 for weights’ calculation). Then the Risk-Adjusted Return Coefficient (RARC) is formulated as follows:
(13) $\quad\gamma_{m,t}=\lambda\bar{\sigma}_{m,t}-(1-\lambda)A_{m,t}$
where $\lambda$ [0,1] is the risk tolerance constant. We defaultly set it as 0.5 to balance risk and return. Section 5.8 demonstrates how we reduce the cubic computational complexity of MPO to a quadratic complexity.
@@
Participate in Global Aggregation. We have evaluated the training RARC $\gamma_{m,t}$ for each cluster. Next, we dynamically incorporate these RARC into the global aggregation process. A higher $\gamma_{m,t}$ indicates that the cluster has experienced a relatively poor local training process. In such scenarios, our objective is to diminish the impact of the current round’s parameter updates on the global aggregation by down-weighting the cluster parameters associated with higher $\gamma_{m,t}$.
To accomplish dynamic aggregation between current round and historical rounds, $FedAdagrad$ employs the following update mechanism: $\bm{\Delta_{m,t}}$ is the difference between the current cluster model and the global model. $\bm{mtn}_{m,t}$ is updated iteratively and collects the historical parameter updates, while $\beta$ balances the historical parameter updates $\bm{mtn}_{m,t-1}$ and $\bm{\Delta}_{m,t}$. A larger $\beta$ indicates the retention of more historical parameter information. $\bm{\upsilon}_{t}$ is an accumulated squared change estimate used to measure changes in parameters and adjust the learning rate to improve the stability and convergence speed of the optimization process:
(14) $\bm{\Delta_{m,t}}=w_{m}(\bm{\Theta}_{m,t}-\bm{\Theta}_{\mathcal{G},t-1})$
$\bm{mtn}_{m,t}=\beta\bm{mtn}_{m,t-1}+(1-\beta)\bm{\Delta_{m,t}},\beta\in[0,1)$
$\bm{\upsilon}_{m,t}=\bm{\upsilon}_{m,t-1}+\bm{\Delta}_{m,t}^{2}$
$\bm{\theta}_{\mathcal{G},t}=\bm{\theta}_{\mathcal{G},t-1}+\eta\frac{\bm{mtn}_{m,t}}{\sqrt{\bm{\upsilon}_{m,t}}+\tau}$
We introduce a dynamic $\beta_{m,t}^{*}$ to replace the static $\beta$ in $FedAdagrad$, which effectively captures the influence of $\gamma_{m,t}$:
(15) $\beta_{m,t}^{*}=\beta+\frac{tanh(ReLU(\gamma_{m,t}))}{\tau}$
$\bm{mtn}_{m,t}=\beta_{m,t}^{*}\bm{mtn}_{m,t-1}+(1-\beta_{m,t}^{*})\bm{\Delta_{m,t}},\beta_{m,t}^{*}\in[0,1)$
@@
When $\gamma_{m,t}>0$, then $\beta_{m,t}^{*}>\beta$, which helps retain more historical parameter updates in the global aggregation. Otherwise, $\beta_{m,t}^{*}=\beta$. Here $\tau$ serves as the smoothness control constant for the range. We suggest setting the temperature to range [10,20], and it depends on the value set for $\beta$.
4.4. Variants of MMiC